OpenStudy (anonymous):

how to find the inverse of my function. F(X)==5x^2+6

OpenStudy (tkhunny):

You cannot find "the" inverse of this function. There are two. You must pick one or report both.

OpenStudy (anonymous):

how do I pick one?

OpenStudy (tkhunny):

x = 5y^2 + 6 x - 6 = 5y^2 (x-6)/5 = y^2 \(f^{-1}(x) = \sqrt{\dfrac{x-6}{5}}\;or\;f^{-1}(x) = -\sqrt{\dfrac{x-6}{5}}\) Generally, find them all (both in this case) and decide what you need.

OpenStudy (abb0t):

Solve for \(x\) in term's of \(y\)

OpenStudy (anonymous):

ooh thats what you do with the power to the 2nd.. you square it!! that's where i got lost. Thanks!!

OpenStudy (tkhunny):

Gotta get your algebra up to speed! You WILL get lost. Good work. Tackle the next one.